3) The vapor pressure of pure ethanol at 60 °C is 0.459 atm. Raoults Law predicts that a solution prepared by dissolving 10. please show all work
Answer #1

Q3. Given : moles of ethanol = 90.0 mmol

moles of naphthalene = 10.0 mmol

Total moles = (moles of ethanol) * (moles of naphthalene)

Total moles = (90.0 mmol) + (10.0 mmol)

Total moles = 100.0 mmol

mole fraction ethanol = (moles of ethanol) / (Total moles)

mole fraction ethanol = (90.0 mmol) / (100.0 mmol)

mole fraction ethanol = 0.900

vapor pressure of solution = (mole fraction ethanol) * (vapor pressure of pure ethanol)

vapor pressure of solution = (0.900) * (0.459 atm)

vapor pressure of solution = 0.413 atm

Q4.

Given : mass NaCl = 58 g

moles NaCl = (mass NaCl) / (molar mass NaCl)

moles NaCl = (58 g) / (58.44 g/mol)

moles NaCl = 0.9925 mol

molality NaCl = (moles NaCl) / (mass of water in kg)

molality NaCl = (0.9925 mol) / (4.01 kg)

molality NaCl = 0.2475 m

increase in boiling point = (i) * (Kb) * (m)

where i = van't Hoff factor = 2 (for NaCl)

Kb = 0.512 oC/m (for water)

m = molality NaCl = 0.2475 m

Substituting the values,

increase in boiling point = (2) * (0.512 oC/m) * (0.2475 m)

increase in boiling point = 0.25 oC

boiling point of solution = (boiling point of pure water) + (increase in boiling point)

boiling point of solution = (100.00 oC) + (0.25 oC)

boiling point of solution = 100.25 oC
Answer #2
  1. correct option is c

  2. bp is 100.25C

answered by: Shivani Sharma